Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 36
Entry time: 26 Apr 2019, 12:28
|
Author: |
Adriano Violante |
Location: |
General |
Severity: |
Information |
Subject: |
Report after dispatch meeting 09.00 |
|
|
SPB/SFX
- IRD commissioning ongoing.
- CRL motors being tested. --> Mechanical issue with a CRL motor is being addressed.
- CSS interlock tests with AE still ongoing.
- AE needs to know when Loop 9 and 10 are expected to run. SPB to check.
- Jungfrau detector: calibration being finalized.
FXE
- Loop 6 is required to drive 2 servo motors. BB will try to bring this into operation this week. NC checked the setup. ==> The loop 9 will be temporarily used. This will be updated either on Monday/Tuesday, or during beamtime. Agreed with FXE.
- Drawers 2 and 3 have to be rebuilt by AE in order to host new ATCA crates. JE needs time plan info from FXE.
SQS
- AE still needs priorities for the next loops. Only Loops 15 and 16 left to be brought up. No plans yet.
- Export of motors config. parameters needed by AE. SQS asks CAS
- Interlocks minor issues being fixed with AE
SCS
- Discussion with EET ongoing concerning the installation of cables for end stations. Space problem. --> This is a long term issue, to be discussed in the other meetings, ongoing.
- Commissioning of the quadrant motion has been completed.
- Loop 12 (chiller) issues with some cables miswiring has been found. AE and EET to check.
- Motors / encoder of the inlinemic component being commissioned.
MID
- For Loop 9 CY is preparing a change request (naming convention for interlocks). AE involved. --> still pending
- Loop 12 will be operable this week --> an 100V input is needed, but not readily available. --> the 100 V were needed to increase the speed of a motor. The alternative solution of using a 48V teminal and a gear box was proposed by AE and accepted by MID.
- Priorities for next loops needed by AE to plan the work. --> Still required.
- Peltier being debugged by MS
- monochromators pico motor in the tunnel is stuck and needs to be fixed. Will be done on Monday.
HED
- Loop9 (Smaract stages) has to be brough up. --> being done. Smaract stages being also tested
- Next loops are 8 followed by the priority-items from loop 6. Loop8 has been started yesterday
- Loop10 and 11 regarding PPL will be needed in June. HED to confirm. --> Feedback from MN required.
- Limit switches (loop8?) are not operable right now due to a 24V terminal missing.
OPT Lasers
- AE support is needed in order to bring the SCS laser up by the end of may. Mail has been sent. NJardon checks. --> Loop extended and running by the end of next week. --> will be done on 06.05.Eplan export will be ready next monday the latest.
XRO
- Motion loop for Hirex 2 has been delayed until June -> update will be done on 11.06.
XPD
- Issue repoeted from CAS: Fast ADC for MCP in SASE3 duplicates data in the 3 % of cases. DAQ is not responsible. Tests have to be made and for that the ADC servers have to undergo to short interruptions. Will be coordinated with XPD. Should not be e problem. CAS to follow the issue up. --> Unclear, CAS not attending.
VAC
DET
AE
ITDM
EET
CAS
SRP
- Schedule changes for the Summer shutdown due to modification of the interlock systems. Additional meetings needed. No info is expected before the beginning of May. --> an uncoordinated email was sent around by MDressel (MPS). SRP will take care of clarifying with MPS what has to be done, when, and will communicate this.
TS
AoB |